OpenStudy (anonymous):

find the vertex, line of symmetry, max/min of f(x) find the vertex, line of symmetry, max/min of f(x) @Mathematics

OpenStudy (anonymous):

f(x)=-2x^2+2x+7

OpenStudy (anonymous):

\[f \prime(x) = -4x + 2\] The root of this equation (also the line of symmetry) is \[x = 1/2\] Plugging this into the original equation yields \[f(x)=-2(1/2)^2+2(1/2)+7=15/2\] Thus the vertex is (1/2,15/2) The range is (-infinity, 15/2)